OpenStudy (anonymous):

Given the function g(x) = 4x − 5, find g(−3).

OpenStudy (callisto):

g(x) = 4x − 5 g(-3) = 4(-3) − 5 = ?

OpenStudy (callisto):

Hint: g(-3) = 4(-3) − 5 = -12 -5 = ... ?

OpenStudy (apoorvk):

plug in '-3' in place of 'x' in '4x-5'

OpenStudy (anonymous):

−17

OpenStudy (callisto):

Yes

OpenStudy (apoorvk):

Yes, right!